OpenStudy (thebookkeeper):

PLEASE HELP! WILL GIVE MEDAL AND FAN! There are 6 girls and 7 boys in a class. A team of 10 players is to be selected from the class. If the selection is random, what is the probability of selecting a team of 4 girls and 6 boys? 0.308 0.367 0.461 0.667

OpenStudy (rational):

probability = (number of favorable outcomes) / (total number of outcomes)

OpenStudy (rational):

find them and take the ratio

OpenStudy (rational):

6+7 = 13 how many ways can you choose 10 players from 13 players ?

OpenStudy (mathmate):

Hint: One way is to pick the students who DON'T make the team, namely one boy and 2 girls. P(BGG)=P(GBG)=P(GGB)=(7/13*6/12*5/11)=35/286 so P(BGG)+P(GBG)+P(GGB) is the required probability. Note: there may be other approaches.